OpenStudy (shamil98):

If you were to place $2,500 in a savings account that pays 5% interest compounded continuously, how much money will you have after 10 years? Assume you make no other deposits or withdrawals. How would I go about finding this?

OpenStudy (snowfire):

You would find what 5% of your account is the first year (which is $125), and add that to the account. Then go on to next year's, and find 5% of your new total, repeating the process.

OpenStudy (luigi0210):

Hmm, you sure Miss Snowfire?

OpenStudy (snowfire):

Not sure if I missed anything, but I wouldn't doubt that fact (I am so clumsy >.>). Maybe I worded it wrong.

OpenStudy (amigatour):

Hang on.. I'm looking through my notes for formulas.

OpenStudy (snowfire):

But using the formula makes this much simpler, P(1+r/n)^nt

OpenStudy (mathstudent55):

\(F = Pe^{rt} \)

OpenStudy (luigi0210):

Dang it, i was writing it out, but I think your formula is the right one, snowfire

OpenStudy (mathstudent55):

When the interest is compounded continuously, you must use the formula with "e."

OpenStudy (snowfire):

Yeah that's what the problem was, I hate e with a passion.

OpenStudy (luigi0210):

Sorry, yea, it's with Pert

OpenStudy (snowfire):

But yeah, P is the starting balance (which is 2500), r is the interest rate as a decimal, and t is the time in years.

OpenStudy (mathstudent55):

You can use the formula with n, as shown above, but you need to take the limit as n approaches infinity. The formula with "e" in it already does that for you.

OpenStudy (snowfire):

Because e is magic, don't question its authority.

OpenStudy (shamil98):

What do those values in the equation mean tho o.o F = Pe^rt

OpenStudy (luigi0210):

Sham, pay attention

OpenStudy (luigi0210):

Snowfire said ir already xD

OpenStudy (luigi0210):

*it

OpenStudy (mathstudent55):

F = Future value P = present value r = annual interest rate written as a decimal t = number of years \(\large F = Pe^{rt} \) \(\large F = 2500e^{0.05 \times 10} \) \( \large F = 2500e^{0.5} \)

OpenStudy (shamil98):

So, now I just evaluate that. F = 2500(2.71)^0.5 ?

OpenStudy (mathstudent55):

yes

OpenStudy (amigatour):

Plug the 2500e^0.05*10 in your calculator.

OpenStudy (shamil98):

4121.8

OpenStudy (amigatour):

That's right.

OpenStudy (mathstudent55):

correct

OpenStudy (shamil98):

thank you all :D